i need to know how to solve this problem:-2(x-4)-4=6
Found 2 solutions by funmath, Nate:
Answer by funmath(2873)

-2(x-4)-4=6
-2(x)-2(-4)-4=6
-2x+8-4=6
-2x+4=6
-2x+4-4=6-4
-2x+0=2
-2x=2
-2x/-2=2/-2
x=-1
:
Check by sugstituting -1 in for x and seeing if both sides =.
-2(-1-4)-4=6
-2(-5)-4=6
10-4=6
6=6 We're right!!!
Happy Calculating!!!

Answer by Nate(3500)

-2(x - 4) - 4 = 6
-2(x - 4) = 10
x - 4 = -5
x = -1